;
DateSerial No.Notes
2018-01-1517884-0190-04+5V power supply failing. Cause of no-boot of electronics as only giving 2.8V until warmed up. Power supply tested on bench ok. Found to be heat sink compound migrated onto output pins over time + some corrosion. Contacts cleaned with servisol + some emery cloth. Re-installed. Works fine.
2018-01-11PMI-1022-1202-31APD heater repaired and tested, refitted to port inner upper position.
2018-01-09PMI-1022-1202-31PCASP-2 poor in flight, looks like APD heater inop. Removed and blanked.
2018-01-04PMI-1022-1202-31PCASP2 calibrated and fitted to port inner upper position
2017-12-15PMI-1022-1202-31PCASP-2 removed after C072 test flight due to absolutely crazy numbers. Inlet jet found to have moved somehow.
2017-12-07PMI-1022-1202-31PCASP-2 sheath flow realigned to the aerosol stream. Cleaned. Inlet jet realigned to laser. Additional cable shield was installed around signal wire from APD to backplane. Wires were shortened and rerouted away from power cables. PCASP-2 was calibrated. PCASP-2 was fitted starboard upper-inner for PICASSO. Quite a strong bin 1 signal when run in the hanger. Vref = 6.7V
2017-10-24PMI-1022-1202-31PCASP-2 removed for C-checks , position blanked.
2017-10-2417884-0190-04PCASP-1 removed for C-checks, position blanked. New dent in PCASP dome?
2017-10-1617884-0190-04PCASP-1 fitted to starboard upper-inner position for ACSIS-2. Tested, Vref=6.16V after warm up.
2017-10-1317884-0190-04PCASP-1 repaired. See PCASP logbook for details of testing. Tip cartridge heaters found to have corrosion which allowed leakage from power lead to chassis. The current limit on the solid state relay was exceeded so this blew (before the rack CBs could pop). Also found that ADP heater was wired into tip heater circuit so only came on when tip heater current was applied. Corrosion was removed and due to some cracks in the ceramic at end of cartridge,the ends were potted up and over the level of the heatshrink. Cartridges were replaced and sealed in with RTV106. Wires leading from heaters to connector were partly wrapped in heatshrink and adhered to the wall of the deacceleration cone to try and reduce turbulence. New Molex 0.062" contacts were crimped onto flying leads. The power supply lines to the APD heater were removed from the tip heater Molex 0.062" connector. A new connector was installed taking power from the input side of the SSR so that power is applied whenever 28V is applied (same as the pump).
2017-09-26PMI-1022-1202-31PCASP-2 fitted for VANAHEIM test flight, portside upper-inner position. Operated and working. Histogram is skewed to small sizes which is a bit odd.
2017-09-2617884-0190-04PCASP-1 fitted for VANAHEIM 2017, portside upper-inner. Tested. 28Vdc current to full-scale when applied tip heaters. Fault has not been fixed. Removed and replaced with PCASP-2.
2017-09-2017884-0190-04PCASP-1 size calibrated in lab with PCASP-2. Desiccant removed and recharged. Tip heater solid state relay defective. Replaced with like-for-like Crydom D1D12 relay.
2017-09-20PMI-1022-1202-31PCASP-2 size calibrated in lab with PCASP-1. Desiccant removed and recharged.
2017-09-15PMI-1022-1202-31PCASP 2 removed from starboard inner upper position for calibration
2017-09-1517884-0190-04PCASP 1 removed from port inner upper position for calibration and repair of non-functional anti-icing heaters
2017-08-02PMI-1022-1202-31PCASP-2 fitted to starboard upper-inner for C027 and CLARIFY.
2017-08-0217884-0190-04PCASP-1 fitted to port upper-inner position for C027 and CLARIFY.
2017-08-02PMI-1022-1202-31PCASP-2 U10 dual op-amp package on multiplexer board very hot. This is due to oscillating during routine operation. Multiplexer board from old PCASP-X fitted in its place. T3 potentiometer adjusted so that Vref similar to with removed board. Desiccant recharged in oven and reinstalled.
2017-07-3117884-0190-04PCASP-1 poor laser ref. 45deg mirror removed and laser OC and mirror cleaned. New o-ring fitted to 45deg mirror as old one flattened. Crystal oscillator removed (usual pitting/marks so blown but not cleaned) and 45deg mirror aligned. Crystal oscillator aligned to maximise laser reference, Vref ~ 7.5V. Desiccant recharged in oven. PCASP-1 calibrated in lab, with PSL and DEHS.
2017-07-18PMI-1022-1202-31PCASP-2 moved from Port inner upper to port upper outer position. Still hot in flight reaching 57C. Gain stage baseline changes when anti ice turned on so flown without.
2017-07-1817884-0190-04PCASP-1 moved from port outer upper to port inner upper position. Shows high bin 1 counts in turns and climbs.
2017-07-11PMI-1022-1202-31PCASP-2 aligned and calibrated in lab, fitted to port upper inner position. Vref ~7.0 v Overheated (+50C) in flight.
2017-07-1017884-0190-04PCASP-1 Fitted port outer upper position. Vref ~7.0 v
2017-06-23PMI-1022-1202-31Removed from Starboard Inner Lower to investigate cyclic noise
2017-06-20PMI-1022-1202-31PCASP-2 displays significant cyclic noise on C014-15. This is most obvious when the heaters are on as (this may be entirely unrelated) the bin 1 counts increases dramatically when the heaters are on. Form of the square wave noise is the same as that seen in VANAHEIM with PCASP-1, period is 2'25" with a positive pulse of 60", but only on bin 1 rather than bins 1-4 on the PCASP-1.
2017-03-24PMI-1022-1202-31Moved from STBD-Upper-Inner to STBD-Lower-Inner
2017-03-1717884-0190-04PCASP-1 tip heater replaced like for like
2017-03-10PMI-1022-1202-31PCASP-2 fitted to stb upper-inner position for Wintex
2017-03-1017884-0190-04Calibrated on bench with PSL and DEHS
2017-03-10PMI-1022-1202-31Calibrated on bench with PSL and DEHS
2017-02-22PMI-1022-1202-31PCASP-2 was removed from stb upper-inner position for MOYA. Blank was attached with shorter non-NAS screws.
2016-12-16PMI-1022-1202-31PCASP-2 fitted to stb upper-inner position for WINTEX17 flying. Vref 7.95v
2016-12-0717884-0190-04PCASP-1 (and PCASP-2) was checked and calibrated in the lab. The calibration included a series of PSLs with a standard threshold table to compare the number concentrations of the two PCASPs.
2016-12-07PMI-1022-1202-31PCASP-2 (and PCASP-1) was checked and calibrated in the lab. The calibration included a series of PSLs with a standard threshold table to compare the number concentrations of the two PCASPs.
2016-12-0617884-0190-04The PCASP-1 sample flow was calibrated with the Gilibrator. Actual sample flow (fa) relationship to PADS flow (f) is; fa = 0.2039f^3 - 0.6888f^2 + 1.5661f - 0.1376
2016-11-0117884-0190-04Removed for c-checks. port-outer-upper
2016-10-1917884-0190-04PCASP1 has been exhibiting a strange cyclic noise pattern in flight, a square-wave with a period of 2.5min and 'on':'off' times of 1:1.5min. During the entire time counts on bins 1-4 are high. All others seem ok. PCASP1 removed from canister in hanger. Gain stage cards removed and inspected for bad contacts. No visual evidence of problem but clean contacts anyway. Reinstall cards and then PCASP. No aircraft power so remains untested.
2016-09-1217884-0190-04PCASP-1 calibrated for diameter and flows (30 Aug 2016). Desiccant baked to renew. Fitted to port side upper-outer position for NAWDEX. Vref = 7.3V
2016-07-2017884-0190-04PCASP-1 removed from aircraft for post-MONSOON clean and calibration.
2016-07-10PMI-1022-1202-31PCASP-2 was removed pre-flight for packing in the hold for the transit back from India.
2016-07-01PMI-1022-1202-31PCASP-2 fitted to starboard lower-inner position for remainder of MONSOON. This was due to PCASPs overheating during low level runs around Lucknow and two the PCASPs were run alternatively in an attempt to alleviate this problem.
2016-04-2917884-0190-04PCASP-1 fitted to port upper-inner position after pre-MONSOON calibration.
2016-02-0817884-0190-04PCASP-1 calibrated in lab. Flow rates also calibrated against Gilibrator. Sheath flow desiccant recharged. PCASP-1 (re)installed port outer-upper position. Vref = 8.4V.
2015-12-1617884-0190-04PCASP1 installed in portside outer-upper position for Jan 2016 test flying. Was externally cleaned and calibrated after ICE-D but may need internal cleaning. Vref = 8.35V.
2015-09-0317884-0190-04PCASP1 removed from port upper-inner after ICE-D. Canister blanked.
2015-07-101PCASP1 fitted to port inner upper.
2015-04-2317884-0190-04PCASP1 installed on portside upper-outer position for spring flying. Some issues in installing the probe in this position. Vref = 9.3V so ~0.5V down on the lab voltages.
2015-04-09PMI-1022-1202-31PCASP1 and 2 simultaneously calibrated in laboratory on 8,9 April. This is the post-COSMICS/CIRCCREX calibration for PCASP2. PCASP2 calibration page updated and calibration files emailed to COSMICS mailing list.
2015-04-0917884-0190-04PCASP1 and 2 simultaneously calibrated in laboratory on 8,9 April. This is the first calibration of the PCASP1 since the repair and realignment at DMT.
2015-04-01PMI-1022-1202-31PCASP2 removed from aircraft for lab calibration.
2015-02-10PMI-1022-1202-31PCASP2 installed portside, inner-upper for CIRCCREX/COSMICS. Tested, Vref=7.1V and ambient air histogram looks reasonable.
2015-02-04PMI-1022-1202-31PCASP2 inlet jet realigned. Probe then calibrated in the lab pre-CIRCCREX/COSMIX. Simultaneous calibration with new MO PCASP-X.
2015-02-02PMI-1022-1202-31Post-PIK&MIX calibration of PCASP2 carried out in lab. Due to work done on PCASP immediately following the lightning strike (9 Dec 2014), the alignment may have changed slightly since the pre-PIK&MIX calibration on 14 Nov 2014.
2015-01-26PMI-1022-1202-31PCASP2 removed from portside, upper-outer canister for alignment and calibration.
2015-01-1317884-0190-04 PCASP1 repaired at DMT. Air filter replaced and RS-422 transceiver chip on DSP module replaced. Cleaned optics and renewed desiccant.
2014-12-12PMI-1022-1202-31PCASP 2 lost coms following a lightening strike during B882 on PIKnMIX at Prestwick. The instrument was removed from the aircraft to conduct fault finding. The fault was traced to the Digital Signal Processing Card. The card was swapped like-for-like with the DSP card from PCASP1, and resolved the fault. PCASP2 was re-fitted to the aircraft.
2014-12-1117884-0190-04Digital Signal Processing card removed from PCASP1. Posted to Prestwick to replace the equivalent card in PCASP2 that had been damaged in lightning strike (the DSP card carries the RS422 serial chips).
2014-11-19PMI-1022-1202-31PCASP2 installed for PIK&MIX, portside upper-outer. This canister is problematic and the installation was quite difficult due to the lip that stands proud of the canister wall halfway down. Turned on to test. There seems to be a change in performance compared to the lab calibration on 14,17th. Vref=7.6V (compared to 8V). The ambient air histogram looks sensible with a possible enhancement of the bin1 counts compared to lab air.
2014-11-17PMI-1022-1202-31PCASP2 calibrated in the lab on 14 and 17 Nov prior to PIK&MIX in Prestwick. Slight misalignment in inlet stream was left uncorrected. Vref=8V.
2014-09-22PMI-1022-1202-31PCASP2 was removed for c-check.
2014-08-27PMI-1022-1202-31Calibrated PCASP2 in lab. Reinstalled port-side upper-outer position.
2014-08-26PMI-1022-1202-31Removed PCASP2 for calibration from port-side upper-outer position.
2014-04-09PMI-1022-1202-31PCASP2 installed portside, outer-upper for GAUGE instrument test flight.
2014-03-20PMI-1022-1202-31PCASP2 calibrated in the the lab. Pre- and post-CAST calibration files emailed to users and PCASP2 calibration webpage updated. Calibration displayed minimal drift between Nov 2013 and Mar 2014.
2014-03-04PMI-1022-1202-31PCASP2 removed for post-CAST calibration.
2013-12-09PMI-1022-1202-31PCASP2 was installed for CAST test flight on port side, upper-outer after lab calibration. Seems to be more communication errors than normal, will check during B820.
2013-12-03PMI-1022-1202-31The PCASP2 was removed from the portside inner-upper position for post-CIRCCREX calibration.
2013-11-0717884-0190-04PCASP1 has not been measuring air flows since the transit back from SALSTICE. It was discovered that the op-amp used as an output buffer for the +10V supply (U10 on Baseline Restoration Module) was electrically damaged (there were no visible signs of damage to package or soldering). A new LT1366 was installed and the PCASP tested in the laboratory. The +10V supply was fine and the measured flows correct.
2013-10-31PMI-1022-1202-31PCASP2 installed for test flight in port side inner-upper position. Tested, Vref = 8.0V.
2013-09-24PMI-1022-1202-31PCASP2 removed from aircraft for C-checks.
2013-06-27PMI-1022-1202-31Calibrated PCASP2 under wing to practice for Exeter.
2013-06-24PMI-1022-1202-31Installed PCASP2 into starboard, upper inner position for COPE test flight.
2013-06-1817884-0190-04Removed PCASP1 for lab calibration
2013-05-0917884-0190-04Re-installed PCASP1 on starboard, inner upper after realignment of inlet stream and calibration.
2013-03-1517884-0190-04Reinstalled PCASP1 onto starboard side, inner-upper after calibration.
2013-03-1517884-0190-04Refitted PCASP1 after calibration. Fitted starboard side, upper-inner.
2013-03-1317884-0190-04Removed PCASP1 for calibration.
2013-03-1117884-0190-04PCASP1 installed after service at DMT. Not calibrated but checked in lab and after installation. Vref=8.5V
2012-12-03PMI-1022-1202-31Installed PCASP 2 into starboard inner-upper in place of PCASP 1. Tested, laser ref voltage non-existant for several minutes until 28V was switched off. Then slowly climbed to 6.2V. Histogram looks very clean for some reason. Blew through 2um glass beads and got a reasonable peak out on the right hand side so finger's crossed it is ok. Voltage is ~1V lower than when in lab so it has been misaligned.
2012-12-0317884-0190-04Removed PCASP 1 from starboard, inner-upper as was not working on B751. Replaced with PCASP 2.
2012-11-2117884-0190-04PCASP 1 installed; starboard inner-upper. Tested 2012-11-23, V=8.5V, ground air histogram looks very good. This is odd as it is this PCASP that was not working at all during the post-SAMBBA cals. So something is perhaps loose on this probe. ADDITIONAL: During B751 this PCASP was not well behaved so there is something wrong with this probe.
2012-10-15PMI-1022-1202-31Removed for C-check.
2012-09-0517884-0190-04Removed PCASP1 and packed into crate for in-hold transport to SAMBBA.
2012-08-31PMI-1022-1202-31Installed PCASP2 onto starboard, inner-upper after calibration. Calibration file: 20120831_P2_cal_results_d.csv
2012-08-3117884-0190-04Calibrated PCASP1 simultaneously with PCASP2. Calibration file: 20120830_P1_cal_results_d.csv Fitted directly beneath PCASP2 in starboard, inner-lower position for direct comparison during test flight, B730.
2012-08-29PMI-1022-1202-31Removed PCASP2 for additional calibration.
2012-08-24PMI-1022-1202-31PCASP2 fitted to starboard inner-upper for test flight, B729. Tested and is working, Vref=6.9V.
2012-07-05PMI-1022-1202-31PCASP#2 installed, starboard upper inner position. Connected to PADS1 computer.
2012-06-29PMI-1022-1202-31Removed PCASP #2 from starboard lower-outer for calibration with PCASP #1.
2012-05-0217884-0190-04Put PCASP1 back onto aircraft. Starboard inner upper position.
2012-04-2617884-0190-04PCASP removed from aircraft and calibrated with PSL and DEHS. No cleaning or aligning was done.
2012-02-2417884-0190-04Installed PCASP #1 on aircraft for test flight. Rewired power-comms patch panel for MEVALI configuration. Tested, Vref = 9.1V
2012-02-2217884-0190-04Calibration of PCASP #1 after overhaul by DMT. An additional cylindrical heater element (or perhaps thermocouple?) had been added by DMT that slid into the brass fitting holding the inlet straw. As such the reassembly of the nose cone has changed. Photographs were taken to add to documentation. Calibration data is in R:DocumentsInstrumentsCloud PhysicsPCASPsCalibrations2012 calibrations20120220_P1_preMEVALI
2012-02-13noneInstrument removed
2012-01-1817884-0190-04PCASP was partially removed from canister to reveal the crystal oscillator. The two thin wires at the base of the unit (not the ones to the detector) go into a Molex connector, one of the wires on the other side had come out of the pin inside the connector (the pins do not crimp very well onto such thin wire). The pin was pushed backwards out of the connector (after depressing the locking flanges). The wire was soldered into a new pin (RS #236-3044), crimped, and pushed back into the connector. The other wire was secure although at some stage should be soldered as well for security. The PCASP was pushed back into the canister and quickly powered-on to test. There is no known cause of this, something must of lightly pulled on the wire while it was on the bench or being moved.
2012-01-1717884-0190-04Refitted. Noticed that wire to oscillator crystal had become detached from Molex connector.
2012-01-1617884-0190-04Realigned inlet stream. Did not do any cleaning or optical realignment due to time constraints. Did a complete PSL calibration of low-gain stage. Did very brief/dodgey calibration of high-gain stage before test flight, P086.
2012-01-0917884-0190-04Removed PCASP #1 for calibration.
2012-01-06nonePost-flight calibration of PCASP for DIAMET after MO realignment of inlet.
2011-12-0717884-0190-04PCASP #1 refitted and basic, on-aircraft calibration carried out with PCL beads. Time: Bead size: 11:08:00-11:14:30 clean DI water 11:16:00-11:20:45 125nm 11:21:15-11:27:45 200nm 11:29:30-11:38:00 1000nm 11:38:30-11:43:00 150nm
2011-12-0617884-0190-04Tubing immediately downstream of air filter had water condensed inside and was kinked. Replaced with a shorter piece of tygon. Filtered air blown through filter for ~2hrs to dry. 0.2um PCL beads used to align inlet. Counts and histogram significantly improved pointing to poor alignment prior to this. One inlet heater cable completely detached. New crimp connectors soldered on and heat shrink used to ensure no shorts. Not a great solution but does the job. Heater current now back to 6A as is expected.
2011-10-03noneRemoved for role change
2011-07-01noneRemoved for BORTAS roll change.
2010-08-272Work done was not entered into the database
2010-08-272Calibrated using DMA for submicron sizes. 2010-08-27
2010-08-272Work done was not entered into the database
2010-08-272Bead cal performed on pan
2010-08-092Baselines adjusted to ensure overlap of all three gain stages. high/mid/low baselines adjusted from 0.0,0.31,0.31 V to 0.0, 0.34, 0.36 V respectively.
2010-08-062Removed for cal.
2010-08-062Optics cleaned and aligned. Leak test performed and passed.
2010-07-092NULL
2010-06-232Removed for calibration and role change
2010-04-222Work done was not entered into the database
2010-04-212PCASP flew on the Dornier at the time of the 146 B522
2010-04-212PCASP flew on the Dornier at the time of the 146 B522
2010-04-192Work done not logged in database
2010-04-092NULL
2010-02-112Refitted after cal
2010-02-102Instrument flow rate increased to 2.5 sccs to get closer to isokinetic sampling. Such a high value is required because we fly at an indicated air speed of 100 m/s which is slower than our true air speed. Optics cleaned and realigned, inlet realigned and cal performed.
2010-02-082Back calibration using DMA performed after optics were realigned while on detachment. No cleaning or realignment was performed as part of the cal
2010-02-072Communications cable broke from its post on the circuit board. Symptoms were no comms from the probe. Reattached and resoldered connection.
2010-02-022NULL
2010-01-272PCASP removed from canister. Repair work took place in the hangar office. Screws on 45 degree mirror optics tightened and aligned to ~8V. Zero check confirmed zero particles and continued good voltage. PCASP was then refitted to the aircraft in same position, and successfully tested on the aircraft..
2010-01-062Fitted for Constrain
2010-01-052Removed for calibration prior to Constrain. Will be transferred by road to Prestwick for re-fitting before flying.
2009-10-292PCASP fitted to PIU.
2009-10-232PCASP calibrated using the DMA and oleic acid.
2009-10-222Laser power supply was clicking for ~30 mins after switch on with the laser flashing for the first few seconds. A new laser power supply was fitted. A number of power supplies were sent back to DMT firstly because the suppied plug was incorrectly wired and this hadnt been spotted but also because the zero load output voltage was measured at 950 V DC not the advertised 1900 VDC. The fitted supply was the third supply and suffered the same low voltage. It was decided that as the laser was functioning at this voltage and the reference voltage was adequate that we would live with this.
2009-10-022Removed for "C" check
2009-09-182Reconnected heater
2009-09-032Fitted after Cal
2009-09-022calibrated using DMA with ammonium sulphate and oleic acid, cleaned and aligned 45 degree mirror laser tip and crystal oscillator, then recalibrated.
2009-08-182remove for cleaning and cal
2009-07-212Full range of available PSL beads sampled by PCASP. In addition modified threshold tables were used to zoom in on the psl spectrum up to 600 nm to give a potentially more accurate cal.
2009-07-172For realignment work.
2009-07-172Cleaned and realigned. Vref = 7.6V
2009-07-172Calibration check using PSLs. Seemed ok.
2009-07-072Calibrated in lab, cleaned and realigned, recalibrated again using DMA.
2009-07-012Removed for calibration
2009-06-052Callibrated with the DMA up to 350 nm using amonium sulphate.
2009-05-282Tested ok with cal beads. 269nm, 1micron good. Flows ok. All power converties/supplies changed.
2009-05-112Quick calibration check before packing.
2009-05-112Removed for transit home.
2009-05-042Calibration checked. OK.
2009-04-192Tested with various sizes of PSLs. Looks ok. Vref approx 8.2V. Electronics very hot (in Muscat) very quickly.